OpenStudy (anonymous):

how do you do 2x+3y=9

OpenStudy (anonymous):

you should ask agen0smith he is a math teacher

OpenStudy (shamil98):

Do what with it? graph it? turn to slope intercept?

OpenStudy (anonymous):

fine slop intercept

OpenStudy (shamil98):

The slope-intercept form is y=mx + b. Solve the equation for y.

OpenStudy (anonymous):

2x+3y=9 try to get y alone 2x+3y=9 subtract 2x from each side 3y = -2x + 9 divide by 3 on both sides y = -2/3x + 3